# EKF–SINDy: Empowering the extended Kalman filter with sparse identification of nonlinear dynamics

### v1

@@ 2.3 Joint estimation of non-autonomous dynamic systems
3:     $\mathbf{F}^{+}(t_{j\!-\!1})=\boldsymbol{\Xi}^{\top}_{f}\frac{\partial\boldsymbol{\Theta}^{\top}_{f}(\boldsymbol{\varkappa})}{\partial\boldsymbol{\varkappa}^{\top}}\bigg{|}_{\hat{\boldsymbol{\varkappa}}^{+}(t_{j\!-\!1})}$
4:     $\hat{\boldsymbol{\varkappa}}^{-}(t_{j})=\hat{\boldsymbol{\varkappa}}^{+}(t_{j\!-\!1})+\Delta t~{}\mathbf{F}^{+}(t_{j\!-\!1})$
@@
6:     $\mathbf{H}^{-}(t_{j})=\boldsymbol{\Xi}^{\top}_{h}\frac{\partial\boldsymbol{\Theta}^{\top}_{h}(\boldsymbol{\varkappa})}{\partial\boldsymbol{\varkappa}^{\top}}\bigg{|}_{\hat{\boldsymbol{\varkappa}}^{-}(t_{j})}$
@@
8:     $\hat{\boldsymbol{\varkappa}}^{+}(t_{j})=\hat{\boldsymbol{\varkappa}}^{-}(t_{j})+\mathbf{G}(t_{j})(\bar{\mathbf{y}}(t_{j})-\boldsymbol{\Xi}^{\top}_{h}\boldsymbol{\Theta}^{\top}_{h}([\hat{\boldsymbol{\varkappa}}^{-}(t_{j}),\mathbf{b}(t_{j})]^{\top}))$
@@
Algorithm 2 EKF-SINDy Joint estimation ($\boldsymbol{\varkappa}=[\mathbf{x},\boldsymbol{\phi}]^{\top}$) of non autonomous dynamical system: online phase.
@@ 3.1 Shear building under seismic excitations
A first numerical case deals with a shear building under seismic excitations. Shear building models are effective in describing the response to lateral excitations, like seismic loads, whenever floors have a sufficiently high out-of-plane stiffness. If this requirement is met, e.g. by way of a minimum slab thickness, the adoption of shear building models in the design phase is allowed by standards such as Eurocode 8 code:EC8 . Here, we have examined the response to seismic excitation of a $2$ storey building assuming regularity in the floor mass and stiffness distribution, such that torsional effects can be neglected and a decoupling of the building response to lateral excitation along the two in-plan directions can be exploited. We thus have employed one degree-of-freedom (dof) per floor to model the response of the building along the horizontal direction. We have assumed the monitoring system to be deployed to record storey displacements $x_{i}(t)$, velocities $\dot{x}_{i}(t)$ and accelerations $\ddot{x}_{i}(t)$, with $i=1,2$, at discrete time steps, obtaining a $6$-dimensional observation vector $\mathbf{y}$. A schematic representation of the shear building model is reported in Fig. 2.
@@
Seismic signals have been taken from the STEAD database art:STEAD . Each seismogram $b(t)$ features the same time duration $T=60$ s and a sampling time step $\Delta t=0.01$ s. The dynamic system is not autonomous due to the presence of this external forcing term. We assume to know the input thanks to the seismic monitoring networks present in many countries, such as in Italy art:Pierleoni18 ; art:DAlessandro19 . On the contrary, we have made the assumption that the storey stiffness is unknown. For instance, when considering reinforced concrete structures, it is challenging to predict how cracking affects the moment of the inertia of column sections, or whether the modelling choices for boundaries have been adequate. Similarly, it is hard to guess how long term degradation effects may have altered the building properties over time. For this reason, we treat the value of the inter-storey stiffness as a stochastic variable featuring a uniform probability density function between $0.5\cdot 10^{6}$ kN///m and $2\cdot 10^{6}$ kN///m. Hence, the stiffness $k$ is the parameter to be here estimated together with the state vector $\mathbf{x}=[x_{1},x_{2},\dot{x}_{1},\dot{x}_{2}]^{\top}$. The second order Eqs. (3.1) can be rewritten at first order using these coordinates, and the dynamics of the system is accordingly described by $n=4$ variables corresponding to the lateral storey displacements and velocities.
To perform the joint estimation of this system, we have constructed a SINDy model for the function $\mathbf{f}$ describing at first order the system, as detailed in Algorithm 1, obtaining $\mathbf{f}\approx\boldsymbol{\Xi}^{\top}_{f}\boldsymbol{\Theta}^{\top}_{f}([\boldsymbol{\varkappa},b]^{\top})$, with $\boldsymbol{\varkappa}=[\mathbf{x},k]^{\top}$. Specifically, the snapshot matrix $\mathbf{X}_{\varkappa b}$ has been constructed by collecting in time the values of the augmented state vector $[\mathbf{x},k]^{\top}$ and for $I=20$ different values of $k$. Polynomial terms up to the second order have been employed in the definition of the function library $\boldsymbol{\Theta}_{f}$, as it can be demonstrated that they can describe the system dynamics, see, e.g. art:RUENG24 . Hence, we have obtained a model for the observation operator $\mathbf{h}$ as in the following
@@
where: $\mathbf{B}_{x}$ and $\mathbf{B}_{\dot{x}}$ are the Boolean matrices that relate displacements and velocities to the observations of the system; $\mathbf{B}_{\ddot{x}}$ is the Boolean matrix that relates the accelerations computed by SINDy to the observed accelerations. The computation of the Jacobian matrices $\mathbf{F}$ and $\mathbf{H}$ has been done as described in Eqs. (15a) and 15b taking advantage of the use of polynomial functions in the library $\boldsymbol{\Theta}_{f}$.
@@
The least square regression used to determine the SINDy weighting coefficients, see Eq. (13), is managed through the python open source package PySINDy art:Kaptanoglu22 . The sparsity promoting regularisation term is set to $L=10^{-2}$. As expected, the adopted second-order library is perfectly suitable to model the dynamics of the system, accounting also for the dependence on $k$.
@@
Hence, we have taken advantage of the combination of SINDy with EKF to perform the joint estimation of shear building whose stiffness is unknown. Specifically, we have taken as initial guess $k=1.01$ kN///m, overestimating by $20\%$ the target value. As previously mentioned, displacements, velocities and accelerations of both floors have been recorded. Although an identifiability assessment has not been carried out, see e.g. art:Chatzis15 , the full observation of the system guarantees the possibility to estimate $k$. In future works, the intention is to explicitly perform an identifiability assessment, for example through the software DAISY art:Bellu07 .
@@
To simulate the outcome of a real monitoring system, signals have been corrupted with white noise featuring a signal to noise ratio equal to $15$. Such level and type of noise is compatible with the use of, e.g., MEMS sensors proc:DAlessandro17 . Before running the online stage of Algorithm 2, we have initialised the state covariance matrix $\mathbf{P}(0)$, the process noise covariance matrix $\mathbf{Q}$, and the observation noise covariance matrix $\mathbf{R}$ as diagonal matrices. The values of the diagonal entries of these matrices are specified in the Appendix. The tuning of these quantities has been carried out through a trial-and-error procedure; however, automatic tuning procedures based for example on genetic algorithms art:Rapp03 or on swarm intelligence art:Laamari15 , can be utilised.
@@
The outcome of the estimation procedure is summarised in Fig. 3 for a seismogram randomly picked from the STEAD database. The mean value of the estimated parameter $\hat{k}$ converges to the target $\bar{k}$ in roughly 10 s reducing in the meanwhile the uncertainty of the estimation, measured by the variance $\sigma_{k}$. The system dynamics is precisely tracked as well, as illustrated in Fig. 4 in terms of a close-up of the kinematic quantities. A confidence interval for the acceleration estimates is not provided, because accelerations are not included in the state vector $\mathbf{x}$. Looking at the displacement and velocity of the first storey, it can be noted that the procedure is strongly tolerant to noise, being the filter estimates $\hat{x}_{1}$ and $\hat{\dot{x}}_{1}$ much closer to the noise-free version of the signals $\hat{\bar{y}}^{x}_{1}$ and $\hat{\bar{y}}^{\dot{x}}_{1}$ than the noise corrupted versions $\bar{y}^{x}_{1}$ and $\bar{y}^{\dot{x}}_{1}$, despite that in the first part of the analysis $k$ is not correctly calibrated yet. Clearly, the matching improved even more after the first $10$ s. The greater discrepancy between acquired recordings and EKF estimates of the accelerations has been due the impossibility of correcting the predicted accelerations, as lateral accelerations are not included in the state vector $\mathbf{x}$.
@@
Figure 3: Shear building. In the top left corner, the shear building model is depicted, with the inter-storey stiffness to be calibrated highlighted in red. In the top right corner, the estimated parameter $\hat{k}_{1}$ is plotted in red against the target $\bar{k}_{1}$ in black. The red shaded area represents $95\%$ confidence interval of the estimates, determined using the posterior covariance. In the other graphs, the evolution of the system response tracked by the filter is plotted in red against the real system dynamics in black. For the sake of presentation, the outcomes of the $60$ s analyses are truncated after $40$ s.
@@
Figure 4: Shear building. Zoom in of Fig. 3: comparison between the evolution in time of the system response tracked by the filter (red dotted lines) and target one (black lines). Red shaded areas represent the $95\%$ confidence intervals determined using the posterior covariance. Black solid lines represent the noise corrupted signals, while black dotted lines their uncorrupted version.
@@
3.2 Partially observed nonlinear dynamic system
A second test case deals with a partially observed nonlinear dynamic system. The system comprises two coupled oscillators characterised by linear stiffness coefficient $k_{i}$, and linear damping coefficient $c_{i}$, with $i=1,2$. These oscillators are linked via linear and quadratic coupling terms controlled by the coefficients $\alpha$ and $\beta$, respectively. The coefficient $\gamma$ dictates instead the contribution of a cubic nonlinearity to the motion of one oscillator. The equations governing the free vibrations of the system are:
@@ 3.2 Partially observed nonlinear dynamic system
for $t=(0,T]$, where: $m_{1}=1$; $m_{2}=1$; $k_{1}=1.0$; $c_{1}=10^{-2}$; $c_{2}=5\cdot 10^{-3}$; $\alpha=-10^{-1}$; $\beta=2\cdot 10^{-3}$; $\gamma=10^{-3}$; $T=200.0$. The time axis is discretised with a sampling time $\Delta T=10^{-2}$. Both the parameters and the time axis have been made dimensionless. Similarly to what considered in Sec. 3.1, the second order Eqs. (3.2) have been rewritten at first order with a simple change of coordinate, ı.e. $(\tilde{z}_{1},\tilde{z}_{2},\tilde{z}_{3},\tilde{z}_{4})=(z_{1},\dot{z}_{1},z_{2},\dot{z}_{2})$, so that the dynamics of the system is described by $n=4$ variables corresponding to the displacements and velocities of the two oscillators. It is worth stressing that, although addressing a real system is beyond the scope of this investigation, several microstructures feature nonlinear interactions between two modes, called internal resonances, that can be effectively modelled as in Eq.(3.2), see e.g. the case of MEMS gyroscope analysed in art:Giorgio21 .
We have assumed to be able to observe only $y(t)=z_{1}(t)$. The aim is therefore to reconstruct the entire system dynamics $\mathbf{f}$ through SINDy by using only snapshots of $z_{1}$, thus considering realistic setting in which both $\mathbf{f}$ and $\mathbf{h}$ must be identified starting only from system observations, with the dimension of the observations smaller than the state of the system, ı.e. $o<n$. We make use of this example to show how to determine a SINDy model when the full state of the system can not be observed, and how to use this model to identify the system parameters. Such parameters can be related to unobserved system components, like the linear stiffness $k_{2}$ of the hidden oscillator and the coupling parameters $\alpha$ and $\beta$. Here, we have employed the time delay embedding technique to enrich the partial observation vector and recover the original dimensionality of the system state. After having trained a SINDy model on the new set of embedded-coordinates, we have proceeded with the system identification with EKF-SINDy.

### v2

@@ 2.3 Joint estimation of non-autonomous dynamical systems
3:     $\mathbf{F}^{+}(t_{j\!-\!1})=\begin{bmatrix}\boldsymbol{\Xi}^{\top}_{f}\frac{\partial\boldsymbol{\Theta}^{\top}_{f}([\boldsymbol{\varkappa},\mathbf{b}]^{\top})}{\partial\mathbf{x}^{\top}}\bigg{|}_{[\hat{\boldsymbol{\varkappa}}^{+},\mathbf{b}](t_{j\!-\!1})}&\boldsymbol{\Xi}^{\top}_{f}\frac{\partial\boldsymbol{\Theta}^{\top}_{f}([\boldsymbol{\varkappa},\mathbf{b}]^{\top})}{\partial\boldsymbol{\phi}^{\top}}\bigg{|}_{[\hat{\boldsymbol{\varkappa}}^{+},\mathbf{b}](t_{j\!-\!1})}\\ \mathbf{0}&\mathbf{0}\end{bmatrix}$
4:     $\hat{\boldsymbol{\varkappa}}^{-}(t_{j})=\hat{\boldsymbol{\varkappa}}^{+}(t_{j\!-\!1})+\Delta t~{}\boldsymbol{\Xi}^{\top}_{f}\boldsymbol{\Theta}^{\top}_{f}([\hat{\boldsymbol{\varkappa}}^{+},\mathbf{b}]^{\top}(t_{j\!-\!1}))$
@@
6:     $\mathbf{H}^{-}(t_{j})=\begin{bmatrix}\boldsymbol{\Xi}^{\top}_{h}\frac{\partial\boldsymbol{\Theta}^{\top}_{h}([\boldsymbol{\varkappa},\mathbf{b}]^{\top})}{\partial\mathbf{x}^{\top}}\bigg{|}_{[\hat{\boldsymbol{\varkappa}}^{-},\mathbf{b}](t_{j})}&\boldsymbol{\Xi}^{\top}_{h}\frac{\partial\boldsymbol{\Theta}^{\top}_{h}([\boldsymbol{\varkappa},\mathbf{b}]^{\top})}{\partial\boldsymbol{\phi}^{\top}}\bigg{|}_{[\hat{\boldsymbol{\varkappa}}^{-},\mathbf{b}](t_{j})}\end{bmatrix}$
@@
8:     $\hat{\boldsymbol{\varkappa}}^{+}(t_{j})=\hat{\boldsymbol{\varkappa}}^{-}(t_{j})+\mathbf{G}(t_{j})(\bar{\mathbf{y}}(t_{j})-\boldsymbol{\Xi}^{\top}_{h}\boldsymbol{\Theta}^{\top}_{h}([\hat{\boldsymbol{\varkappa}}^{-},\mathbf{b}]^{\top}(t_{j})))$
@@
Algorithm 2 EKF-SINDy joint estimation ($\boldsymbol{\varkappa}=[\mathbf{x},\boldsymbol{\phi}]^{\top}$) of non-autonomous dynamical systems: online phase
@@ 3.1 Shear building under seismic excitations
A first numerical case deals with a shear building under seismic excitations. Shear building models are effective in describing the response to lateral excitations, like seismic loads, whenever floors have a sufficiently high out-of-plane stiffness. If this requirement is met, e.g. by way of a minimum slab thickness, the adoption of shear building models in the design phase is allowed by standards such as Eurocode 8 [51]. Here, we have examined the response to seismic excitation of a $2$ storey building assuming regularity in the floor mass and stiffness distribution, such that torsional effects can be neglected and a decoupling of the building response to lateral excitation along the two in-plan directions can be exploited. We thus have employed one degree-of-freedom (dof) per floor to model the response of the building along the horizontal direction. We have assumed the monitoring system to be deployed to record storey displacements $x_{i}(t)$, velocities $\dot{x}_{i}(t)$ and accelerations $\ddot{x}_{i}(t)$, with $i=1,2$, at discrete time steps, obtaining a $6$-dimensional observation vector $\mathbf{y}$. A schematic representation of the shear building model is reported in Fig. 3.
@@
Seismic signals have been taken from the STEAD database [52]. Each seismogram $b(t)$ features the same time duration $60$ s and a sampling time step of $0.01$ s. Here, we have subsampled each signal getting a final sampling of $0.001$ s to avoid stability issue with the Euler forward integration. The dynamical system is not autonomous due to the presence of this external forcing term. We assume to know the input thanks to the seismic monitoring networks present in many countries, such as in Italy [53, 46]. On the contrary, we have made the assumption that the storey stiffness is unknown. For instance, when considering reinforced concrete structures, it is challenging to predict how cracking affects the moment of the inertia of column sections, or whether the modelling choices for boundaries have been adequate. Similarly, it is hard to guess how long term degradation effects may have altered the building properties over time. For this reason, we treat the value of the inter-storey stiffness as a stochastic variable featuring a uniform probability density function between $0.5\cdot 10^{6}$ kN///m and $2\cdot 10^{6}$ kN///m. Hence, the stiffness $k$ is the parameter to be here estimated together with the state vector $\mathbf{x}=[x_{1},x_{2},\dot{x}_{1},\dot{x}_{2}]^{\top}$. The second order Eq. (22) can be rewritten at first order using these coordinates, and the dynamics of the system is accordingly described by $n=4$ variables corresponding to the lateral storey displacements and velocities.
To perform the joint estimation of this system, we have constructed a SINDy model for the function $\mathbf{f}$ describing at first order the system, as detailed in Algorithm 1, obtaining $\mathbf{f}\approx\boldsymbol{\Xi}^{\top}_{f}\boldsymbol{\Theta}^{\top}_{f}([\boldsymbol{\varkappa},b]^{\top})$, with $\boldsymbol{\varkappa}=[\mathbf{x},k]^{\top}$. Specifically, the snapshot matrix $\mathbf{X}_{\varkappa b}$ has been constructed by collecting in time the values of the augmented state vector $[\mathbf{x},k]^{\top}$ and for $I=20$ different values of $k$. Polynomial terms up to the second order have been employed in the definition of the function library $\boldsymbol{\Theta}_{f}$, as it can be demonstrated that they can describe the system dynamics, see, e.g. [26]. Hence, we have obtained a model for the observation operator $\mathbf{h}$ as in the following
@@
where: $\mathbf{B}_{x}$ and $\mathbf{B}_{\dot{x}}$ are the Boolean matrices that relate displacements and velocities to the observations of the system; $\mathbf{B}_{\ddot{x}}$ is the Boolean matrix that relates the accelerations computed by SINDy to the observed accelerations. The computation of the Jacobian matrices $\mathbf{F}$ and $\mathbf{H}$ has been done as described in Eqs. (15a) and (15b) taking advantage of the use of polynomial functions in the library $\boldsymbol{\Theta}_{f}$.
@@
The least square regression used to determine the SINDy weighting coefficients, see Eq. (13), is managed through the python open source package PySINDy [54]. Sparsity is promoted using STLSQ with threshold $L=10^{-2}$ and regularisation strength $\delta_{r}=0.05$. The value for $\delta_{r}$ is the default setting in the PySINDy package. In the considered case studies, we have found that adjusting $L$ alone has been sufficient to control the sparsity of the system. As expected, the adopted library of second-order polynomials is perfectly suitable to model the dynamics of the system, accounting also for the dependence on $k$.
@@
Hence, we have taken advantage of the combination of SINDy with EKF to perform the joint estimation of shear building whose stiffness is unknown. Specifically, we have taken as initial guess $k=1.01\cdot 10^{6}$ kN///m, overestimating by $20\%$ the target value. As previously mentioned, displacements, velocities and accelerations of both floors have been recorded. Although an identifiability assessment has not been carried out, see e.g. [55], the full observation of the system guarantees the possibility to estimate $k$. In future works, the intention is to explicitly perform an identifiability assessment, for example through the software DAISY [56].
@@
To simulate the outcome of a real monitoring system, signals have been corrupted with white noise featuring a signal to noise ratio equal to $15$. Such level and type of noise is compatible with the use of, e.g., MEMS sensors [57]. Before running the online stage of Algorithm 2, we have initialised the state covariance matrix $\mathbf{P}(0)$, the process noise covariance matrix $\mathbf{Q}$, and the observation noise covariance matrix $\mathbf{R}$ as diagonal matrices. The values of the diagonal entries of these matrices are specified in Appendix B. The tuning of these quantities has been carried out through a trial-and-error procedure (a practical guidance to expedite this procedure is included in Appendix B as well); however, automatic tuning procedures based on Bayesian Optimisation with Gaussian Processes [58], on genetic algorithms [59] or on swarm intelligence [60], can be utilised.
@@
The outcome of the estimation procedure is summarised in Fig. 4 for a seismogram randomly picked from the STEAD database. The mean value of the estimated parameter $\hat{k}$ converges to the target $\bar{k}$ in roughly 20 s reducing the uncertainty of the estimation, measured by the standard deviation $\sigma_{k}$ determined using the posterior covariance computed by the EKF. The system dynamics is precisely tracked as well, as illustrated in Fig. 5 in terms of a close-up of the kinematic quantities. A confidence interval for the acceleration estimates is not provided, because accelerations are not included in the state vector $\mathbf{x}$. The procedure is strongly tolerant to noise, being the filter estimates (for example $\hat{x}_{1}$) much closer to the noise-free version of the signals ($\hat{\bar{y}}^{x}_{1}$) than the noise corrupted versions ($\bar{y}^{x}_{1}$), despite that in the first part of the analysis $k$ is not correctly calibrated yet. Clearly, the matching improved even more after the first $20$ s.
@@
Figure 4: Shear building. In the top graph, the estimated parameter $\hat{k}$ is plotted in red against the target $\bar{k}$ in black. The red shaded area represents $95\%$ confidence interval of the estimates, determined using the posterior covariance. In the other graphs, the evolution of the system response tracked by the filter is plotted in red against the real system dynamics in black. For the sake of presentation, the outcomes of the $60$ s analyses are truncated after $40$ s.
@@
Figure 5: Shear building. Zoom in of Fig. 4: comparison between the evolution in time of the system response tracked by the filter (red dotted lines) and target one (black lines) for incorrect (on the left) and updated (on the right) value of $k$. Red shaded areas represent the $95\%$ confidence intervals determined using the posterior covariance. Black solid lines represent the noise corrupted signals, while black dotted lines their uncorrupted version.
@@
3.2 Partially observed nonlinear dynamical system
A second test case deals with a partially observed nonlinear dynamical system. The system comprises two coupled oscillators characterised by linear stiffness coefficient $k_{i}$, and linear damping coefficient $c_{i}$, with $i=1,2$. These oscillators are linked via linear and quadratic coupling terms controlled by the coefficients $\alpha$ and $\beta$, respectively. The coefficient $\gamma$ dictates instead the contribution of a cubic nonlinearity to the motion of one oscillator. The equations governing the free vibrations of the system are:
@@ 3.2 Partially observed nonlinear dynamical system
where: $m_{1}=1$; $m_{2}=1$; $k_{1}=1.0$; $c_{1}=2\cdot 10^{-2}$; $c_{2}=1.95\cdot 10^{-2}$; $\alpha=-10^{-1}$; $\beta=2\cdot 10^{-3}$; $\gamma=10^{-3}$. The time axis, of length $200.0$, is discretised with a time step of $\Delta t=10^{-2}$. Both the parameters and the time axis have been made dimensionless. Similarly to what considered in Sec. 3.1, the second order Eqs. (24) have been rewritten at first order with a simple change of coordinate, ı.e. $(\tilde{z}_{1},\tilde{z}_{2},\tilde{z}_{3},\tilde{z}_{4})=(z_{1},\dot{z}_{1},z_{2},\dot{z}_{2})$, so that the dynamics of the system is described by $n=4$ variables corresponding to the displacements and velocities of the two oscillators. It is worth stressing that, although addressing a real system is beyond the scope of this investigation, several microstructures feature nonlinear interactions between two modes, called internal resonances, that can be effectively modelled as in Eq.(24), see e.g. the case of MEMS gyroscope analysed in [61].
We have assumed to be able to observe only $y(t)=z_{1}(t)$. The aim is therefore to reconstruct the entire system dynamics $\mathbf{f}$ through SINDy by using only snapshots of $z_{1}$, thus considering realistic setting in which both $\mathbf{f}$ and $\mathbf{h}$ must be identified starting only from system observations, with the dimension of the observations smaller than the state of the system, ı.e. $o<n$. We make use of this example to show how to determine a SINDy model when the full state of the system can not be observed, and how to use this model to identify the system parameters. Such parameters can be related to unobserved system components, like the linear stiffness $k_{2}$ of the hidden oscillator and the coupling parameters $\alpha$ and $\beta$. Here, we have employed the time delay embedding technique, presented in Sec. 2.2.2, to enrich the partial observation vector and recover the original dimensionality of the system state.
We have verified that without time delay embedding, it is impossible to get a SINDy model that accurately predicts the system dynamics $\mathbf{f}$, thereby allowing system identification. In this scenario, the library $\boldsymbol{\Theta}_{f}$ would be solely a function of $z_{1}$ and the system parameters to be estimated (as we will highlight in the next subsections, we have separately addressed the estimation of $k_{2}$ and the coupling parameters). A SINDy model based on such a function library cannot catch some fundamental aspects of the system dynamics, such as the dependence of the damping force on $\dot{z}_{1}$. Considering that it is always possible to recover $\dot{z_{1}}$ from $z_{1}$, we have also attempted to train a SINDy model by using both $z_{1}$ and $\dot{z_{1}}$, but even in this case, the prediction accuracy of the model is insufficient for system identification purposes. This inadequacy persists regardless the threshold hyperparameter $L$, as it is fundamentally linked to the impossibility of including the dependence on the second oscillator. These analyses have confirmed the necessity of using time delay coordinates. After having trained a SINDy model on the new set of embedded-coordinates, we have thus proceeded with the system identification with EKF-SINDy.